OpenStudy (anonymous):

Find f^-1(x) and state any restrictions on the domain of f^-1(x)

OpenStudy (anonymous):

\[f(x)=\frac{ x }{ x-2 }\] \[x \neq 2 \]

OpenStudy (anonymous):

you need the inverse?

OpenStudy (anonymous):

Yes. I know I need to switch f(x) and x, but I dunno how you would simplify from there

OpenStudy (anonymous):

you can take \[y=\frac{x}{x-2}\] switch \(x\) and \(y\) to write \[x=\frac{y}{y-2}\] and solve for \(y\)

OpenStudy (anonymous):

you get \[x=\frac{y}{y-2}\]\[x(y-2)=y\] as a first step then multiply out

OpenStudy (anonymous):

\[xy-2x=y\] put the \(y\)'s on one side of the equal sign, the \(2x\) on the other \[xy-y=2x\]

OpenStudy (anonymous):

factor out the \(y\) on the left \[(x-1)y=2x\]

OpenStudy (anonymous):

Okay. I've got it from there

OpenStudy (anonymous):

I didn't think to subtract y over ^^;;

OpenStudy (anonymous):

and finally divide the domain will be all real numbers except \(x=1\) which is not surprising because that was the range of \(f\)

OpenStudy (anonymous):

answer should be \[y=\frac{2x}{x-1}\]

OpenStudy (anonymous):

yeah you got to get the \(y\)' all on one side good luck

OpenStudy (anonymous):

Thanks.